OpenStudy (callisto):

Tutor question #3 Simplify: \[\sqrt{5-\sqrt{21}}\] (conditions: pen, paper, no calculator, done in 2 minutes) *

OpenStudy (lgbasallote):

what if x = \(\sqrt{5-\sqrt{21}}\) then \(x^2 = 5 - \sqrt{21}\) \[x^2 - 5 = -\sqrt{21}\] \[(x^2 - 5)^2 = 21\] oh goodness quadratic o.O

OpenStudy (anonymous):

actually it is a 4th degree equation

OpenStudy (callisto):

\[(x^2-5)^2=21\]\[x^4-10x^2+25=21\]\[x^4-10x^2+4=0\]\[x^4-10x^2+4=0\]\[x^2=\frac{10\pm \sqrt{(-10)^6-4(4)}}{2}\]\[x^2=5\pm \sqrt{21}\]Sorry... are you sure that it is 'simplified'?

OpenStudy (callisto):

Oh.that's power 2, typo

OpenStudy (anonymous):

apparently it is also \[\sqrt{\frac{7}{2}}-\sqrt{\frac{3}{2}}\] i remember seeing something like this before but i am not sure i remember how to go from one to the other

OpenStudy (callisto):

I need to know how to work that out.....

OpenStudy (anonymous):

cant seem to get it, i think it was a trick

OpenStudy (callisto):

............................ It... was ... a ... question ... asked ... when ... my friend applied for a summer job ...

OpenStudy (anonymous):

really? doing what??

OpenStudy (callisto):

tutor - teaching high school students

OpenStudy (anonymous):

hmm i wonder what answer they wanted

OpenStudy (callisto):

Same here...

OpenStudy (callisto):

When I was doing some exercises few days ago, I saw similar questions, but clearer, like this: express \(\sqrt{28-2\sqrt{147}}\) in the form of \(\sqrt{x}-\sqrt{y}\). I can still handle this. But that one, I failed :(

OpenStudy (anonymous):

multiplying by the conjugate give \(\frac{2}{\sqrt{5+\sqrt{21}}}\) think

myininaya (myininaya):

any examples anywhere?

OpenStudy (callisto):

example?!

myininaya (myininaya):

You want to show \[\sqrt{5-\sqrt{21}} \text{ equals } \frac{\sqrt{7}-\sqrt{3}}{2} ?\] was just wondering if you have an example for writing that one thing in that other form

myininaya (myininaya):

oops sqrt(2) on bottom

myininaya (myininaya):

how did you get that identity?

OpenStudy (callisto):

From experience...

myininaya (myininaya):

lol great experience I want to prove that identity lol

OpenStudy (callisto):

First, \((\sqrt{a} - \sqrt{b})^2\) = a + b -2\sqrt{ab} For a>b \[\sqrt{5-\sqrt{21}} = \sqrt{5-2\sqrt{\frac{21}{4}}}\] Now, a+b = 5 => a=5-b ab = 21/4 (5-b)b = 21/4 -4b^2 + 20b - 21 =0 b=1.5 or b =3.5 (rejected) a = 5 - 1.5 = 3.5 So, it is \(\sqrt{\frac{7}{2}}-\sqrt{\frac{3}{2}}\) Does that make sense?

myininaya (myininaya):

ok i see that identity :)

myininaya (myininaya):

that one is easy to prove

myininaya (myininaya):

maybe because it is an actual identity, right? :p

OpenStudy (callisto):

Perfect square is perfect :)

OpenStudy (callisto):

Does that make sense? Apart from the latex fail...

myininaya (myininaya):

very interesting i wouldn't have thought of that

OpenStudy (callisto):

I'm going to post the link to satellite73's post then. He doesn't even come to check..

myininaya (myininaya):

Great work @Callisto :)

OpenStudy (callisto):

And thank you for all your time!!!!